\begin{thebibliography}{67}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Clarkson et~al.(2023)Clarkson, Ubaru, and Yang]{clarkson2023capacity}
Kenneth~L Clarkson, Shashanka Ubaru, and Elizabeth Yang.
\newblock Capacity analysis of vector symbolic architectures.
\newblock \emph{arXiv:2301.10352}, pages 1--49, 2023.

\bibitem[Diao et~al.(2021)Diao, Kleyko, Rabaey, and
  Olshausen]{diao2021generalized}
Cameron Diao, Denis Kleyko, Jan~M Rabaey, and Bruno~A Olshausen.
\newblock Generalized learning vector quantization for classification in
  randomized neural networks and hyperdimensional computing.
\newblock In \emph{International Joint Conference on Neural Networks (IJCNN)},
  pages 1--9, 2021.

\bibitem[Dua and Graff(2017)]{Dua:2019}
Dheeru Dua and Casey Graff.
\newblock {UCI} machine learning repository, 2017.
\newblock URL \url{http://archive.ics.uci.edu/ml}.

\bibitem[Emruli et~al.(2013)Emruli, Gayler, and Sandin]{EmruliAnalogical2013}
Blerim Emruli, Ross~W Gayler, and Fredrik Sandin.
\newblock Analogical mapping and inference with binary spatter codes and sparse
  distributed memory.
\newblock In \emph{International Joint Conference on Neural Networks (IJCNN)},
  pages 1--8, 2013.

\bibitem[Fern{\'a}ndez-Delgado et~al.(2014)Fern{\'a}ndez-Delgado, Cernadas,
  Barro, and Amorim]{fernandez2014we}
Manuel Fern{\'a}ndez-Delgado, Eva Cernadas, Sen{\'e}n Barro, and Dinani Amorim.
\newblock Do we need hundreds of classifiers to solve real world classification
  problems?
\newblock \emph{The Journal of Machine Learning Research}, 15\penalty0
  (1):\penalty0 3133--3181, 2014.

\bibitem[Fodor and Pylyshyn(1988)]{fodor1988connectionism}
Jerry~A Fodor and Zenon~W Pylyshyn.
\newblock Connectionism and cognitive architecture: A critical analysis.
\newblock \emph{Cognition}, 28\penalty0 (1-2):\penalty0 3--71, 1988.

\bibitem[Frady et~al.(2018)Frady, Kleyko, and Sommer]{FradyCapacity2018}
E~Paxon Frady, Denis Kleyko, and Friedrich~T Sommer.
\newblock A theory of sequence indexing and working memory in recurrent neural
  networks.
\newblock \emph{Neural Computation}, 30\penalty0 (6):\penalty0 1449--1513,
  2018.

\bibitem[Frady et~al.(2020)Frady, Kent, Olshausen, and
  Sommer]{frady2020resonator}
E~Paxon Frady, Spencer~J Kent, Bruno~A Olshausen, and Friedrich~T Sommer.
\newblock Resonator networks, 1: An efficient solution for factoring
  high-dimensional, distributed representations of data structures.
\newblock \emph{Neural Computation}, 32\penalty0 (12):\penalty0 2311--2331,
  2020.

\bibitem[Frady et~al.(2021)Frady, Kleyko, Kymn, Olshausen, and
  Sommer]{fradyfunctions2021}
E~Paxon Frady, Denis Kleyko, Christopher~J Kymn, Bruno~A Olshausen, and
  Friedrich~T Sommer.
\newblock Computing on functions using randomized vector representations.
\newblock \emph{arXiv:2109.03429}, pages 1--33, 2021.

\bibitem[Frady et~al.(2022)Frady, Kleyko, Kymn, Olshausen, and
  Sommer]{frady2022computing}
E~Paxon Frady, Denis Kleyko, Christopher~J Kymn, Bruno~A Olshausen, and
  Friedrich~T Sommer.
\newblock Computing on functions using randomized vector representations (in
  brief).
\newblock In \emph{{Neuro-Inspired Computational Elements Conference (NICE)}},
  pages 115--122, 2022.

\bibitem[Frady et~al.(2023)Frady, Kleyko, and Sommer]{FradySDR2021}
E~Paxon Frady, Denis Kleyko, and Friedrich~T. Sommer.
\newblock Variable binding for sparse distributed representations: Theory and
  applications.
\newblock \emph{{IEEE Transactions on Neural Networks and Learning Systems}},
  34\penalty0 (5):\penalty0 2191--2204, 2023.

\bibitem[Ganesan et~al.(2021)Ganesan, Gao, Gandhi, Raff, Oates, Holt, and
  McLean]{GanesanLearning2021}
Ashwinkumar Ganesan, Hang Gao, Sunil Gandhi, Edward Raff, Tim Oates, James
  Holt, and Mark McLean.
\newblock Learning with holographic reduced representations.
\newblock In \emph{{Advances in Neural Information Processing Systems
  (NeurIPS)}}, pages 1--15, 2021.

\bibitem[Gayler(1998)]{GaylerMAP1998}
Ross~W Gayler.
\newblock Multiplicative binding, representation operators \& analogy.
\newblock In \emph{Advances in Analogy Research: Integration of Theory and Data
  from the Cognitive, Computational, and Neural Sciences}, pages 1--4, 1998.

\bibitem[Gayler(2003)]{gayler2003vector}
Ross~W Gayler.
\newblock Vector symbolic architectures answer {J}ackendoff's challenges for
  cognitive neuroscience.
\newblock In \emph{Joint International Conference on Cognitive Science
  (ICCS/ASCS)}, pages 133--138, 2003.

\bibitem[Gayler and Levy(2009)]{GaylerIsomorphism2009}
Ross~W Gayler and Simon~D Levy.
\newblock A distributed basis for analogical mapping: New frontiers in analogy
  research.
\newblock In \emph{{New frontiers in Analogy Research, Second International
  Conference on the Analogy (ANALOGY)}}, pages 165--174, 2009.

\bibitem[Ge and Parhi(2020)]{ge2020classification}
Lulu Ge and Keshab~K Parhi.
\newblock Classification using hyperdimensional computing: A review.
\newblock \emph{{IEEE Circuits and Systems Magazine}}, 20\penalty0
  (2):\penalty0 30--47, 2020.

\bibitem[Gosmann and Eliasmith(2019)]{gosmann2019vector}
Jan Gosmann and Chris Eliasmith.
\newblock Vector-derived transformation binding: An improved binding operation
  for deep symbol-like processing in neural networks.
\newblock \emph{Neural computation}, 31\penalty0 (5):\penalty0 849--869, 2019.

\bibitem[Graves et~al.(2014)Graves, Wayne, and Danihelka]{graves2014neural}
Alex Graves, Greg Wayne, and Ivo Danihelka.
\newblock Neural {T}uring machines.
\newblock \emph{arXiv:1410.5401}, pages 1--25, 2014.

\bibitem[Greff et~al.(2020)Greff, Van~Steenkiste, and
  Schmidhuber]{greff2020binding}
Klaus Greff, Sjoerd Van~Steenkiste, and J{\"u}rgen Schmidhuber.
\newblock On the binding problem in artificial neural networks.
\newblock \emph{arXiv:2012.05208}, 2020.

\bibitem[Heddes et~al.(2022)Heddes, Nunes, Givargis, Nicolau, and
  Veidenbaum]{heddes2022hyperdimensional}
Mike Heddes, Igor Nunes, Tony Givargis, Alexandru Nicolau, and Alex Veidenbaum.
\newblock Hyperdimensional hashing: A robust and efficient dynamic hash table.
\newblock In \emph{{ACM/IEEE Design Automation Conference (DAC)}}, pages
  907--912, 2022.

\bibitem[Hernandez-Cane et~al.(2021)Hernandez-Cane, Matsumoto, Ping, and
  Imani]{hernandez2021onlinehd}
Alejandro Hernandez-Cane, Namiko Matsumoto, Eric Ping, and Mohsen Imani.
\newblock {OnlineHD}: Robust, efficient, and single-pass online learning using
  hyperdimensional system.
\newblock In \emph{Design, Automation \& Test in Europe Conference \&
  Exhibition (DATE)}, pages 56--61, 2021.

\bibitem[Hersche et~al.(2022)Hersche, Karunaratne, Cherubini, Benini,
  Sebastian, and Rahimi]{HerscheContinualLearn2022}
Michael Hersche, Geethan Karunaratne, Giovanni Cherubini, Luca Benini, Abu
  Sebastian, and Abbas Rahimi.
\newblock Constrained few-shot class-incremental learning.
\newblock In \emph{{Conference on Computer Vision and Pattern Recognition
  (CVPR)}}, pages 9057--9067, 2022.

\bibitem[Hersche et~al.(2023)Hersche, Zeqiri, Benini, Sebastian, and
  Rahimi]{hersche2023neuro}
Michael Hersche, Mustafa Zeqiri, Luca Benini, Abu Sebastian, and Abbas Rahimi.
\newblock A neuro-vector-symbolic architecture for solving {R}aven's
  progressive matrices.
\newblock \emph{Nature Machine Intelligence}, 5\penalty0 (4):\penalty0
  363--375, 2023.

\bibitem[Hopfield(1982)]{hopfield1982neural}
John~J Hopfield.
\newblock Neural networks and physical systems with emergent collective
  computational abilities.
\newblock \emph{Proceedings of the national academy of sciences}, 79\penalty0
  (8):\penalty0 2554--2558, 1982.

\bibitem[Imani et~al.(2017)Imani, Kong, Rahimi, and Rosing]{imani2017voicehd}
Mohsen Imani, Deqian Kong, Abbas Rahimi, and Tajana Rosing.
\newblock {VoiceHD}: Hyperdimensional computing for efficient speech
  recognition.
\newblock In \emph{International Conference on Rebooting Computing (ICRC)},
  pages 1--8, 2017.

\bibitem[Kanerva(1988)]{kanerva1988sparse}
Pentti Kanerva.
\newblock \emph{Sparse distributed memory}.
\newblock MIT press, 1988.

\bibitem[Kanerva(1997)]{kanerva1997fully}
Pentti Kanerva.
\newblock Fully distributed representation.
\newblock In \emph{Real World Computing Symposium (RWC)}, pages 358--365, 1997.

\bibitem[Kanerva(2000)]{KanervaLearning2000}
Pentti Kanerva.
\newblock Large patterns make great symbols: An example of learning from
  example.
\newblock In \emph{International Workshop on Hybrid Neural Systems}, volume
  1778 of \emph{Lecture Notes in Computer Science}, pages 194--203, 2000.

\bibitem[Kanerva(2009)]{kanerva2009hyperdimensional}
Pentti Kanerva.
\newblock Hyperdimensional computing: An introduction to computing in
  distributed representation with high-dimensional random vectors.
\newblock \emph{Cognitive Computation}, 1\penalty0 (2):\penalty0 139--159,
  2009.

\bibitem[Kang et~al.(2022)Kang, Khaleghi, Rosing, and Kim]{kang2022openhd}
Jaeyoung Kang, Behnam Khaleghi, Tajana Rosing, and Yeseong Kim.
\newblock {OpenHD}: A {GPU}-powered framework for hyperdimensional computing.
\newblock \emph{{IEEE Transactions on Computers}}, 71\penalty0 (11):\penalty0
  2753--2765, 2022.

\bibitem[Kleyko et~al.(2015)Kleyko, Osipov, Gayler, Khan, and
  Dyer]{KleykoBees2015}
Denis Kleyko, Evgeny Osipov, Ross~W Gayler, Asad~I Khan, and Adrian~G Dyer.
\newblock Imitation of honey bees' concept learning processes using vector
  symbolic architectures.
\newblock \emph{Biologically Inspired Cognitive Architectures}, 14:\penalty0
  57--72, 2015.

\bibitem[Kleyko et~al.(2021)Kleyko, Kheffache, Frady, Wiklund, and
  Osipov]{kleyko2020density}
Denis Kleyko, Mansour Kheffache, E~Paxon Frady, Urban Wiklund, and Evgeny
  Osipov.
\newblock Density encoding enables resource-efficient randomly connected neural
  networks.
\newblock \emph{IEEE Transactions on Neural Networks and Learning Systems},
  32\penalty0 (8):\penalty0 3777--3783, 2021.

\bibitem[Kleyko et~al.(2022{\natexlab{a}})Kleyko, Bybee, Kymn, Olshausen,
  Khosrowshahi, Nikonov, Sommer, and Frady]{kleyko2022integer}
Denis Kleyko, Connor Bybee, Christopher~J Kymn, Bruno~A Olshausen, Amir
  Khosrowshahi, Dmitri~E Nikonov, Friedrich~T Sommer, and E~Paxon Frady.
\newblock Integer factorization with compositional distributed representations.
\newblock In \emph{{Neuro-Inspired Computational Elements Conference (NICE)}},
  pages 73--80, 2022{\natexlab{a}}.

\bibitem[Kleyko et~al.(2022{\natexlab{b}})Kleyko, Davies, Frady, Kanerva, Kent,
  Olshausen, Osipov, Rabaey, Rachkovskij, Rahimi, and Sommer]{kleyko2021vector}
Denis Kleyko, Mike Davies, Edward~Paxon Frady, Pentti Kanerva, Spencer~J. Kent,
  Bruno~A. Olshausen, Evgeny Osipov, Jan~M. Rabaey, Dmitri~A. Rachkovskij,
  Abbas Rahimi, and Friedrich~T. Sommer.
\newblock Vector symbolic architectures as a computing framework for emerging
  hardware.
\newblock \emph{Proceedings of the IEEE}, 110\penalty0 (10):\penalty0
  1538--1571, 2022{\natexlab{b}}.

\bibitem[Kleyko et~al.(2022{\natexlab{c}})Kleyko, Rachkovskij, Osipov, and
  Rahimi]{kleyko2021survey1}
Denis Kleyko, Dmitri~A Rachkovskij, Evgeny Osipov, and Abbas Rahimi.
\newblock A survey on hyperdimensional computing aka vector symbolic
  architectures, part {I}: Models and data transformations.
\newblock \emph{{ACM Computing Surveys}}, 55\penalty0 (6):\penalty0 1--40,
  2022{\natexlab{c}}.

\bibitem[Kleyko et~al.(2023)Kleyko, Rachkovskij, Osipov, and
  Rahimi]{kleyko2021survey2}
Denis Kleyko, Dmitri~A Rachkovskij, Evgeny Osipov, and Abbas Rahimi.
\newblock A survey on hyperdimensional computing aka vector symbolic
  architectures, part {II}: Applications, cognitive models, and challenges.
\newblock \emph{{ACM Computing Surveys}}, 55\penalty0 (9):\penalty0 1--52,
  2023.

\bibitem[Krotov and Hopfield(2016)]{krotov2016dense}
Dmitry Krotov and John~J Hopfield.
\newblock Dense associative memory for pattern recognition.
\newblock \emph{Advances in neural information processing systems}, 29, 2016.

\bibitem[Kussul et~al.(1991)Kussul, Rachkovskij, and
  Baidyk]{kussul1991associative}
Ernst~M Kussul, Dmitri~A Rachkovskij, and Tatyana~N Baidyk.
\newblock Associative-projective neural networks: Architecture, implementation,
  applications.
\newblock In \emph{International Conference on Neural Networks and Their
  Applications (NEURO)}, pages 463--476, 1991.

\bibitem[Laiho et~al.(2015)Laiho, Poikonen, Kanerva, and
  Lehtonen]{LaihoSparse2015}
Mika Laiho, Jussi~H Poikonen, Pentti Kanerva, and Eero Lehtonen.
\newblock High-dimensional computing with sparse vectors.
\newblock In \emph{{IEEE Biomedical Circuits and Systems Conference (BioCAS)}},
  pages 1--4, 2015.

\bibitem[Langenegger et~al.(2023)Langenegger, Karunaratne, Hersche, Benini,
  Sebastian, and Rahimi]{langenegger2023memory}
Jovin Langenegger, Geethan Karunaratne, Michael Hersche, Luca Benini, Abu
  Sebastian, and Abbas Rahimi.
\newblock In-memory factorization of holographic perceptual representations.
\newblock \emph{{Nature Nanotechnology}}, 18\penalty0 (5):\penalty0 479--485,
  2023.

\bibitem[Neubert et~al.(2019)Neubert, Schubert, and
  Protzel]{NeubertRobotics2019}
Peer Neubert, Stefan Schubert, and Peter Protzel.
\newblock An introduction to hyperdimensional computing for robotics.
\newblock \emph{KI - K{\"u}nstliche Intelligenz}, 33\penalty0 (4):\penalty0
  319--330, 2019.

\bibitem[Nunes et~al.(2022)Nunes, Heddes, Givargis, Nicolau, and
  Veidenbaum]{nunes2022graphhd}
Igor Nunes, Mike Heddes, Tony Givargis, Alexandru Nicolau, and Alex Veidenbaum.
\newblock {GraphHD}: Efficient graph classification using hyperdimensional
  computing.
\newblock In \emph{Design, Automation \& Test in Europe Conference \&
  Exhibition (DATE)}, pages 1485--1490, 2022.

\bibitem[Nunes et~al.(2023)Nunes, Heddes, Givargis, and
  Nicolau]{nunes2022extension}
Igor Nunes, Mike Heddes, Tony Givargis, and Alexandru Nicolau.
\newblock An extension to basis-hypervectors for learning from circular data in
  hyperdimensional computing.
\newblock In \emph{{ACM/IEEE Design Automation Conference (DAC)}}, 2023.

\bibitem[Osipov et~al.(2017)Osipov, Kleyko, and Legalov]{osipov2017associative}
Evgeny Osipov, Denis Kleyko, and Alexander Legalov.
\newblock Associative synthesis of finite state automata model of a controlled
  object with hyperdimensional computing.
\newblock In \emph{Annual Conference of the IEEE Industrial Electronics Society
  (IECON)}, pages 3276--3281, 2017.

\bibitem[Paszke et~al.(2017)Paszke, Gross, Chintala, Chanan, Yang, DeVito, Lin,
  Desmaison, Antiga, and Lerer]{paszke2017automatic}
Adam Paszke, Sam Gross, Soumith Chintala, Gregory Chanan, Edward Yang, Zachary
  DeVito, Zeming Lin, Alban Desmaison, Luca Antiga, and Adam Lerer.
\newblock Automatic differentiation in {PyT}orch.
\newblock In \emph{Advances in Neural Information Processing Systems
  (NeurIPS)}, pages 1--4, 2017.

\bibitem[Paszke et~al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan,
  Killeen, Lin, Gimelshein, Antiga, et~al.]{paszke2019pytorch}
Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory
  Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, et~al.
\newblock {PyT}orch: An imperative style, high-performance deep learning
  library.
\newblock In \emph{Advances in Neural Information Processing Systems
  (NeurIPS)}, 2019.

\bibitem[Plate(1991)]{PlateAlgebra1991}
Tony~A Plate.
\newblock Holographic reduced representations: Convolution algebra for
  compositional distributed representations.
\newblock In \emph{International Joint Conference on Artificial Intelligence
  (IJCAI)}, pages 30--35, 1991.

\bibitem[Plate(1994)]{PlateAnalogical1994}
Tony~A Plate.
\newblock Estimating analogical similarity by dot-products of holographic
  reduced representations.
\newblock In \emph{Advances in Neural Information Processing Systems (NIPS)},
  pages 1109--1116, 1994.

\bibitem[Plate(1995)]{plate1995holographic}
Tony~A Plate.
\newblock Holographic reduced representations.
\newblock \emph{IEEE Transactions on Neural Networks}, 6\penalty0 (3):\penalty0
  623--641, 1995.

\bibitem[Rachkovskij and Slipchenko(2012)]{RachkovskijAnalogy2012}
Dmitri~A Rachkovskij and Serge~V Slipchenko.
\newblock Similarity-based retrieval with structure-sensitive sparse binary
  distributed representations.
\newblock \emph{Computational Intelligence}, 28\penalty0 (1):\penalty0
  106--129, 2012.

\bibitem[Rachkovskij et~al.(2005)Rachkovskij, Slipchenko, Kussul, and
  Baidyk]{rachkovskiy2005sparse}
Dmitri~A Rachkovskij, Sergey~V Slipchenko, Ernst~M Kussul, and Tatyana~N
  Baidyk.
\newblock Sparse binary distributed encoding of scalars.
\newblock \emph{Journal of Automation and Information Sciences}, 37\penalty0
  (6):\penalty0 12--23, 2005.

\bibitem[Rahimi et~al.(2016{\natexlab{a}})Rahimi, Benatti, Kanerva, Benini, and
  Rabaey]{rahimi2016hyperdimensional}
Abbas Rahimi, Simone Benatti, Pentti Kanerva, Luca Benini, and Jan~M Rabaey.
\newblock Hyperdimensional biosignal processing: A case study for {EMG-based}
  hand gesture recognition.
\newblock In \emph{International Conference on Rebooting Computing (ICRC)},
  pages 1--8, 2016{\natexlab{a}}.

\bibitem[Rahimi et~al.(2016{\natexlab{b}})Rahimi, Kanerva, and
  Rabaey]{rahimi2016robust}
Abbas Rahimi, Pentti Kanerva, and Jan~M Rabaey.
\newblock A robust and energy-efficient classifier using brain-inspired
  hyperdimensional computing.
\newblock In \emph{International Symposium on Low Power Electronics and Design
  (ISLPED)}, pages 64--69, 2016{\natexlab{b}}.

\bibitem[Rahimi et~al.(2019)Rahimi, Kanerva, Benini, and
  Rabaey]{RahimiBiosignal2019}
Abbas Rahimi, Pentti Kanerva, Luca Benini, and Jan~M Rabaey.
\newblock Efficient biosignal processing using hyperdimensional computing:
  Network templates for combined learning and classification of {ExG} signals.
\newblock \emph{Proceedings of the IEEE}, 107\penalty0 (1):\penalty0 123--143,
  2019.

\bibitem[Rahimi and Recht(2007)]{rahimi2007random}
Ali Rahimi and Benjamin Recht.
\newblock Random features for large-scale kernel machines.
\newblock In \emph{Advances in Neural Information Processing Systems (NIPS)},
  volume~20, pages 1--8, 2007.

\bibitem[Ramsauer et~al.(2020)Ramsauer, Sch{\"a}fl, Lehner, Seidl, Widrich,
  Adler, Gruber, Holzleitner, Pavlovi{\'c}, Sandve,
  et~al.]{ramsauer2020hopfield}
Hubert Ramsauer, Bernhard Sch{\"a}fl, Johannes Lehner, Philipp Seidl, Michael
  Widrich, Thomas Adler, Lukas Gruber, Markus Holzleitner, Milena Pavlovi{\'c},
  Geir~Kjetil Sandve, et~al.
\newblock Hopfield networks is all you need.
\newblock \emph{arXiv preprint arXiv:2008.02217}, 2020.

\bibitem[Rasmussen and Eliasmith(2014)]{RasmussenSpiking2014}
Daniel Rasmussen and Chris Eliasmith.
\newblock A spiking neural model applied to the study of human performance and
  cognitive decline on {R}aven's advanced progressive matrices.
\newblock \emph{Intelligence}, 42:\penalty0 53--82, 2014.

\bibitem[Renner et~al.(2022)Renner, Supic, Danielescu, Indiveri, Olshausen,
  Sandamirskaya, Sommer, and Frady]{renner2022neuromorphic}
Alpha Renner, Lazar Supic, Andreea Danielescu, Giacomo Indiveri, Bruno~A
  Olshausen, Yulia Sandamirskaya, Friedrich~T Sommer, and E~Paxon Frady.
\newblock Neuromorphic visual scene understanding with resonator networks.
\newblock \emph{arXiv:2208.12880}, pages 1--15, 2022.

\bibitem[Schlegel et~al.(2022)Schlegel, Neubert, and
  Protzel]{schlegel2022comparison}
Kenny Schlegel, Peer Neubert, and Peter Protzel.
\newblock A comparison of vector symbolic architectures.
\newblock \emph{Artificial Intelligence Review}, 55\penalty0 (6):\penalty0
  4523--4555, 2022.

\bibitem[Simon et~al.(2022)Simon, Pale, Teijeiro, and
  Atienza]{simon2022hdtorch}
William~Andrew Simon, Una Pale, Tomas Teijeiro, and David Atienza.
\newblock {HDTorch}: Accelerating hyperdimensional computing with {GP-GPU}s for
  design space exploration.
\newblock In \emph{IEEE/ACM International Conference on Computer-Aided Design
  (ICCAD)}, pages 1--8, 2022.

\bibitem[Smolensky(1990)]{SmolenskyTensor1990}
Paul Smolensky.
\newblock Tensor product variable binding and the representation of symbolic
  structures in connectionist systems.
\newblock \emph{Artificial Intelligence}, 46:\penalty0 159--216, 1990.

\bibitem[Smolensky et~al.(2022)Smolensky, McCoy, Fernandez, Goldrick, and
  Gao]{smolensky2022neurocompositional}
Paul Smolensky, Richard McCoy, Roland Fernandez, Matthew Goldrick, and Jianfeng
  Gao.
\newblock Neurocompositional computing: From the central paradox of cognition
  to a new generation of {AI} systems.
\newblock \emph{AI Magazine}, 43\penalty0 (3):\penalty0 308--322, 2022.

\bibitem[Teeters et~al.(2022)Teeters, Kleyko, Kanerva, and
  Olshausen]{teeters2022separating}
Jeffrey~L Teeters, Denis Kleyko, Pentti Kanerva, and Bruno~A Olshausen.
\newblock On separating long-and short-term memories in hyperdimensional
  computing.
\newblock \emph{Frontiers in Neuroscience}, 16:\penalty0 1--19, 2022.

\bibitem[Thomas et~al.(2021)Thomas, Dasgupta, and
  Rosing]{thomas2021theoretical}
Anthony Thomas, Sanjoy Dasgupta, and Tajana Rosing.
\newblock Theoretical foundations of hyperdimensional computing.
\newblock \emph{Journal of Artificial Intelligence Research (JAIR)},
  72:\penalty0 215--249, 2021.

\bibitem[Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez,
  Kaiser, and Polosukhin]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin.
\newblock Attention is all you need.
\newblock \emph{Advances in neural information processing systems}, 30, 2017.

\bibitem[von~der Malsburg(1986)]{MalsburgAssemblies1986}
Christoph von~der Malsburg.
\newblock Am {I} thinking assemblies?
\newblock In \emph{Brain Theory}, pages 161--176, 1986.

\bibitem[Yerxa et~al.(2018)Yerxa, Anderson, and
  Weiss]{yerxa2018hyperdimensional}
Thomas Yerxa, Alexander Anderson, and Eric Weiss.
\newblock The hyperdimensional stack machine.
\newblock \emph{Cognitive Computing}, pages 1--2, 2018.

\end{thebibliography}

\pdfoutput=1 

\documentclass[twoside,11pt]{article}

\usepackage{blindtext}

\usepackage[preprint]{jmlr2e}

\usepackage{amsmath}
\usepackage{booktabs}
\usepackage{pgf}
\usepackage{import}
\usepackage{tikz,tkz-euclide,pdftexcmds,calc}
\usetikzlibrary{arrows,shapes.geometric}
\usetikzlibrary{positioning}
\usetikzlibrary{automata}
\usetikzlibrary{arrows,shapes,calc}
\usepackage{circuitikz}
\usepackage{algorithmic}
\usepackage[linesnumbered,ruled,noend]{algorithm2e}
\usepackage{bbm}
\usepackage{multirow}

\usepackage{amssymb}
\usepackage{pifont}
\usepackage{bm} 
\usepackage{makecell} 
\definecolor{ao(english)}{rgb}{0.0, 0.5, 0.0}
\newcommand{\cmark}{\color{ao(english)} \ding{51}}%
\newcommand{\xmark}{\color{red} \ding{55}}%

\hypersetup{hidelinks}

\newcommand{\libraryname}{Torchhd}

\usepackage{lastpage}
\jmlrheading{24}{2023}{1-\pageref{LastPage}}{3/23}{6/23}{23-0300}{Mike Heddes, Igor Nunes, Pere Vergés, Denis Kleyko, Danny Abraham, Tony Givargis, Alexandru Nicolau, and Alexander Veidenbaum}
\ShortHeadings{\libraryname{}}{Heddes, Nunes, Vergés, Kleyko, Abraham, Givargis, Nicolau, and Veidenbaum}

\firstpageno{1}

\begin{document}

\title{\libraryname{}: An Open Source Python Library to Support Research on Hyperdimensional~Computing and Vector~Symbolic~Architectures}

\author{\name Mike Heddes\textsuperscript{\normalfont{1}} \email mheddes@uci.edu
    \AND
    \name Igor Nunes\textsuperscript{\normalfont{1}} \email igord@uci.edu
    \AND
    \name Pere Vergés\textsuperscript{\normalfont{1}} \email pvergesb@uci.edu
    \AND
    \name Denis Kleyko\textsuperscript{\normalfont{2}} \email denis.kleyko@ri.se
    \AND
    \name Danny Abraham\textsuperscript{\normalfont{1}} \email dannya1@uci.edu
    \AND
    \name Tony Givargis\textsuperscript{\normalfont{1}} \email givargis@uci.edu
    \AND
    \name Alexandru Nicolau\textsuperscript{\normalfont{1}} \email nicolau@ics.uci.edu
    \AND
    \name Alexander Veidenbaum\textsuperscript{\normalfont{1}} \email alexv@ics.uci.edu
    \AND
   \textsuperscript{\normalfont{1}} \addr Department of Computer Science, University of California, Irvine, CA 92617, USA
   \AND
   \textsuperscript{\normalfont{2}} \addr
   Intelligent Systems Lab, Research Institutes of Sweden, 164 40 Kista, Sweden
}

\editor{Sebastian Schelter}

\maketitle

\begin{abstract}
Hyperdimensional computing (HD), also known as vector symbolic architectures (VSA), is a framework for computing with distributed representations by exploiting properties of random high-dimensional vector spaces. 
The commitment of the scientific community to aggregate and disseminate research in this particularly multidisciplinary area has been fundamental for its advancement. 
Joining these efforts, we present \libraryname{}, a high-performance open source Python library for HD/VSA. 
\libraryname{} seeks to make HD/VSA more accessible and serves as an efficient foundation for further research and application development.
The easy-to-use library builds on top of PyTorch and features state-of-the-art HD/VSA functionality, clear documentation, and implementation examples from well-known publications. Comparing publicly available code with their corresponding \libraryname{} implementation shows that experiments can run up to 100$\times$ faster. \libraryname{} is available at:
\url{https://github.com/hyperdimensional-computing/torchhd}.
\end{abstract}

\begin{keywords}
  hyperdimensional computing, vector symbolic architectures, distributed representations, machine learning, symbolic AI, Python library
\end{keywords}

\section{Introduction}
\label{sec:introduction}

Hyperdimensional computing (HD)~\citep{kanerva2009hyperdimensional}, also known as vector symbolic architectures (VSA)~\citep{gayler2003vector}, is a computing framework capable of forming compositional distributed representations. 
It originated within artificial intelligence and cognitive science~\citep{SmolenskyTensor1990, kussul1991associative, PlateAlgebra1991} from attempts to address limitations of earlier distributed representations~\citep{MalsburgAssemblies1986,fodor1988connectionism}.  
HD/VSA forms a ``concept space'' by exploiting the geometry and algebra of high-dimensional spaces. 
The central idea is to represent information with randomly generated vectors, called \textit{hypervectors}.
Together with a set of operations on these hypervectors, HD/VSA can represent compositional structures, which, in turn, enables features such as reasoning by analogy~\citep{PlateAnalogical1994, KanervaLearning2000, GaylerIsomorphism2009, RachkovskijAnalogy2012, KleykoBees2015} and cognitive computing~\citep{EmruliAnalogical2013, RasmussenSpiking2014, hersche2023neuro}.   

While the foundational ideas of HD/VSA have been around for some time, it was only recently that it started gaining momentum both theoretically~\citep{FradyCapacity2018, thomas2021theoretical,clarkson2023capacity} and practically, attracting attention from the wider machine learning community~\citep{graves2014neural, RahimiBiosignal2019, NeubertRobotics2019, GanesanLearning2021, HerscheContinualLearn2022}.  
This interest is partially driven by the observation that the strengths of HD/VSA complement some of the limitations of current artificial neural networks~\citep{greff2020binding,smolensky2022neurocompositional}, leading to hybrid approaches that could be viewed as neuro-symbolic~\citep{hersche2023neuro}.

In face of the growing interest, there have been substantial efforts to consolidate and disseminate HD/VSA by its research community. 
The challenge for these initiatives has been the fact that HD/VSA involves many disciplines, including machine learning, neuroscience, electrical engineering, artificial intelligence, mathematics, and cognitive science.
Keeping up with the latest advances is not trivial as publications are spread across different venues and disciplines. 
It is, therefore, important to pursue consolidation and integration efforts, which will facilitate further research endeavors. 
Despite detailed surveys of various aspects of the area~\citep{ge2020classification, schlegel2022comparison, kleyko2021survey1, kleyko2021survey2}, much less attention, however, has been paid towards streamlining the translation of concepts, methodologies, and algorithms into a coherent open source software.

To address this issue, we developed \libraryname{}---an open source library for HD/VSA. 
Up to date with the latest advances in the area, \libraryname{} aims to lower the barrier of entry to HD/VSA for novices and provides a high-performance execution platform for experienced researchers.
To support a wide variety of HD/VSA primitives, applications, and research directions; the provided abstractions are designed to be modular, allowing users to adopt the aspects that support their workflow.
Our design philosophy focuses first on ease-of-use, aiming to be accessible without limiting expressiveness.
The ease of use allows for application-driven research to focus on the conceptual and methodological aspects of the project while, at the same time, \libraryname{} enables the study of individual elements of HD/VSA for theoretical research.
The second goal of the library is to provide high-performance execution, enabling much faster evaluation of novel algorithms without hurting the Python idiomatic development experience. 
We believe that \libraryname{} will benefit the HD/VSA and the broader machine learning communities based on its following features:
1) state-of-the-art HD/VSA methods for transforming data into hypervector;
2) support for automatic differentiation for research on hybrid neuro-symbolic models;
3) Python idiomatic interface design for a flexible, developer-first experience;
4) high-performance execution, which means that applications can run orders of magnitude faster;
5) interfaces for data sets commonly used in the literature to evaluate and benchmark HD/VSA methods (129 data sets currently).

\section{\libraryname}
\label{sec:library}

\libraryname{} builds on PyTorch~\citep{paszke2019pytorch}, a library for high-performance tensor computation. 
PyTorch and its ecosystem provide many machine learning utilities, support for various hardware accelerators, and built-in automatic differentiation (autodiff) \citep{paszke2017automatic}. 
Building on PyTorch enables the exchange of knowledge between HD/VSA and the wider machine learning community, and facilitates the cross-fertilization of ideas. 

\libraryname{} supports six common HD/VSA models, namely binary spatter codes (BSC) \citep{kanerva1997fully}, multiply-add-permute (MAP) \citep{GaylerMAP1998}, holographic reduced representations (HRR), Fourier HRR (FHRR) \citep{plate1995holographic}, sparse block codes (SBC) \citep{LaihoSparse2015,FradySDR2021}, and vector-derived transformation binding (VTB) \citep{gosmann2019vector}. 
These models differ in their specification of the high-dimensional space and their implementations of the fundamental operations: superposition and binding of hypervectors.
Support for (F)HRR enabled the implementation of vector function architectures \citep{fradyfunctions2021, frady2022computing} which extends an approach for implementing large-scale kernel machines via random Fourier features \citep{rahimi2007random}.
Our implementations are verified with extensive automatic unit testing. 
We also place emphasis on clear documentation, which is available at: 
\url{https://torchhd.readthedocs.io}

The library has six modules which provide the following functionality:
\vspace{-0.5em}%
\begin{itemize}
    \setlength\itemsep{-0.1em}
    
    \item \texttt{functional}: functions for generating hypervectors \citep{rachkovskiy2005sparse, nunes2022extension}; the operations on hypervectors \citep{schlegel2022comparison}; and a resonator network for factorizing hypervectors \citep{frady2020resonator, renner2022neuromorphic, kleyko2022integer, langenegger2023memory}.
    
    \item \texttt{embeddings}: classes for transforming scalars or feature vectors into hypervectors, ranging from simple hypervector lookups till similarity-preserving transformations of feature vectors \citep{kleyko2020density,thomas2021theoretical}, some of which approximate well-known kernels~\citep{rahimi2007random, frady2022computing}.  
    
    \item \texttt{models}: common classification models such as the centroid model with class prototypes and its various training algorithms \citep{rahimi2016hyperdimensional, imani2017voicehd, hernandez2021onlinehd, nunes2022graphhd}, learning vector quantization \citep{diao2021generalized}, and regularized least squares  \citep{kleyko2020density}.
    
    \item \texttt{memory}: methods for long-term storage and retrieval of hypervectors, such as sparse distributed memory \citep{kanerva1988sparse}, inspired by the cerebellum \citep{teeters2022separating}, the (modern) Hopfield network \citep{hopfield1982neural, krotov2016dense}, and the attention mechanism \citep{vaswani2017attention, ramsauer2020hopfield}.
    
    \item \texttt{structures}: classes for HD/VSA data structures \citep{kleyko2021vector} such as hash tables, graphs, binary trees, and finite state automata \citep{osipov2017associative, yerxa2018hyperdimensional, heddes2022hyperdimensional}. This module facilitates the development of classical algorithms using HD/VSA. 
    
    \item \texttt{datasets}: convenient access to 126 classification and 3 regression data sets commonly used in the literature. This includes a classification benchmark of UCI machine learning repository data sets \citep{Dua:2019} created by \citet{fernandez2014we}. All data sets are interoperable with the PyTorch ecosystem.
\end{itemize}

\begin{table}[tb]
    \centering
    \footnotesize
    \setlength{\tabcolsep}{5pt}
    \begin{tabular}{l|cccccccccc}
         \toprule
         Library & BSC & MAP & HRR & FHRR & \makecell{Memory\\models} & \makecell{Data\\structures} & \makecell{Resonator\\ Network} & \makecell{Auto-\\diff} & Data sets \\
         \midrule
         OpenHD & \xmark & \cmark & \xmark & \xmark & \xmark & \xmark & \xmark & \xmark & \xmark\\
         HDTorch & \cmark & \cmark & \xmark & \xmark & \xmark & \xmark & \xmark & $\bm{\pm}$ & \xmark\\
         VSA Toolbox & \cmark & \cmark & \cmark & \cmark & \xmark & \xmark & \xmark & \xmark & \xmark \\
         \libraryname{} (ours) & \cmark & \cmark & \cmark & \cmark & \cmark & \cmark & \cmark & \cmark & \cmark \\ 
         \bottomrule
    \end{tabular}
    \caption{\label{tab:lib-compare}A qualitative assessment of the functionality of HD/VSA software.}
\end{table}

We note that several other software for HD/VSA exist.
Table~\ref{tab:lib-compare} contrasts the features of \libraryname{} with those of OpenHD \citep{kang2022openhd}, HDTorch \citep{simon2022hdtorch}, and VSA Toolbox \citep{schlegel2022comparison}. 
As part of their review on HD/VSA models, \cite{schlegel2022comparison} implement a broader range of HD/VSA models, however, their VSA Toolbox misses many of the features that make \libraryname{} particularly useful for research and development of HD/VSA, see, for example, Table~\ref{tab:lib-compare}.
Importantly, \libraryname{} is the first general software library for HD/VSA since the other software were developed for more specific purposes.

To demonstrate the performance of \libraryname{}, we implemented three classification tasks and compared the execution time to the original publicly available code. 
All the experiments ran on a machine with 20 Intel Xeon Silver 4114 CPUs, 93~GB of RAM, and 4 Nvidia TITAN Xp GPUs, only a single CPU or GPU was used. 
The results in Table~\ref{tab:performance} show that \libraryname{} is faster for all examples with an average of 24$\times$ and 54$\times$ faster for CPU and GPU, respectively.
The provided abstractions ensure performant code by avoiding overly sequential execution, reinforced by our support for batch processing.

\begin{table}[h]
  \centering
  \footnotesize
  \begin{tabular}{l|ccc}
    \toprule
    Paper  & Original & CPU & GPU\\
    \midrule
    EU languages~\citep{rahimi2016robust} & 13111.73 & 542.83 (24$\times$) & \textbf{125.89} (104$\times$)\\
    EMG gestures~\citep{rahimi2016hyperdimensional} & 1152.32 & \textbf{28.44} (41$\times$) & 28.95 (40$\times$)\\
    VoiceHD~\citep{imani2017voicehd} & 277.97 & 37.74 (7$\times$) & \textbf{16.17} (17$\times$)\\
    \bottomrule
\end{tabular}
  \caption{\label{tab:performance} Execution time in seconds (and speedup)}
\end{table}
\section{Conclusions and Future Work}
\label{sec:conclusion}

\libraryname{} is a Python library for hyperdimensional computing (HD) a.k.a. vector symbolic architectures (VSA) that aims to be simple, versatile, and highly performant.
The library builds on PyTorch, thus, enabling the exchange of knowledge between the HD/VSA and the broader machine learning communities as well as making PyTorch's vast ecosystem of deep learning tools available for HD/VSA. 
\libraryname{} is designed to include any HD/VSA model and already supports binary spatter codes, multiply-add-permute, (Fourier) holographic reduced representations, sparse block codes, and vector-derived transformation binding. 
We plan to keep the library up-to-date with novel developments in the area by including more HD/VSA models and supporting a broader spectrum of learning algorithms, such as differentiable learning, regression, density estimation, and clustering. 
Our work joins the community’s vital effort to consolidate and disseminate HD/VSA research.

\newpage
\acks{%
We would like to thank Ross W. Gayler for helpful discussions on the design choices and desired functionality of the library;
Rishikanth Chandrasekaran, Dheyay Desai, Jenny Lee, and Xiaofan Yu for their contributions to the library; and finally, Anthony Thomas for his support while resolving an issue with the random projection implementation.
\newline
DK received funding from the European Union's Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 839179.
}

\vskip 0.2in
\bibliography{references}

\end{document}